OpenStudy (anonymous):

Is my answer correct? f(x) = 8x + 2 and g(x) = 5x - 3. What is the value of f/g (4)? I got the answer 2.

OpenStudy (anonymous):

Or is the answer 8?

OpenStudy (anonymous):

Yippee! You're getting the hang of it.

OpenStudy (anonymous):

Wait so it is 2?

OpenStudy (anonymous):

Go with your first!

OpenStudy (anonymous):

So 2?

OpenStudy (anonymous):

34/17 = 2

OpenStudy (anonymous):

I know but after that there is a 4 so do I multiply them together? 2x4 = 8

OpenStudy (anonymous):

\[\frac{ 8*4+2 }{ 5*4-3 }=\frac{ 32+2 }{ 20-3 }=\frac{ 34 }{ 17 }\] The 4 indicates the value to be substituted in for x.

OpenStudy (anonymous):

Ok thank you I understand this now!!!